Sin78°-sin18°+sin30°-sin42°. Value of this trigonometry from chapter 5

Grade:11

1 Answers

gaurav
7 years ago
sin78-sin18+sin30-sin42=sin78-sin42-sin18+sin30=2cos(120/2).sin18-sin18+sin30=2cos60.sin18-sin18+sin30=(2×1/2).sin18-sin18+sin30=sin18-sin18+1/2=1/2